Mrs John is 5 times older than her son .3 years ago the product of their ages was 185 how old is Mrs John
1
Expert's answer
2020-07-29T11:41:28-0400

Mrs J - > X years ago,

Her son - > Y years old 

"X =5*Y\\"

"\begin{alignedat}{2}

 (5&Y-3)&(Y-3) = 185
\end{alignedat}"

"5*Y^{2}-18*Y-176=0\\"


a=5, b=-18, c=-176=0


"D=b^{2}-4*a*c,"

"D=(-18)^{2}-4*1*(-176)=62^2,"


"Y\ufeff1=(-b+\\sqrt{D})\/(2*a),\\\\\nY\ufeff1=(-(-18)+\\sqrt{62^2})\/(2*5)=8,\\\\\nY\ufeff2=(-b-\\sqrt{D})\/(2*a), \\\\\nY2=(-(-18)-\\sqrt{62^2})\/(2*5)=-4.4. \\\\"


The age must be greater than 0, so Y=8,

X=5*Y=5*8=40.

Mrs John is 40 years old, her son is 8 years old,

Checking: (40-3)*(8-3)=185.

Answer: Now Mrs John is 40 years old, her son is 8 years old.